# Novel Lower Bounds on $M/G/k$ Scheduling

### v2

@@ 4.1. Proof overview
For any job size distribution $S$ and arrival rate $\lambda$, the expected total work in the $M/G/k$ system under any scheduling policy is lower bounded by the expected total work in the ISQ-$k$ system.
We prove Theorem 4.3 using a sample path coupling argument in Section 6. In fact, we prove a more general lower bound under an arbitrary arrival sequence.
@@
Using Theorem 4.3, we prove in Theorems 6.2 and 6.3 that the expected relevant work of the Sep-ISQ-$k$ system and the AR-ISQ-$k$ system each lower bound the $M/G/k$ expected relevant work under any scheduling policy for all arrival rates $\lambda$.
@@
Now that we’ve proven that Sep-ISQ-$k$ and AR-ISQ-$k$ lower bound mean relevant work in the M/G/$k$, our remaining goal is to characterize the expected relevant work of the Sep-ISQ-$k$ and AR-ISQ-$k$ systems. As an intermediate result, a key step is to compute the expected total work of the ISQ-$k$ system. We analyze the expected total work of the ISQ-$k$ system using the DiffeDrift method, see Section 8.1.
@@
Here $v_{1}$ and $u_{1}$ are functions defined as the solutions of differential equations in Definitions 9.1 and 9.2. Equations 4.5 and 4.6 correspond to Propositions 7.3 and 9.3. These results are proved in Sections 7 and 9.
Sep-ISQ-$k$ and AR-ISQ-$k$ relevant work
@@
Now, it remains to analyze the mean relevant work in Sep-ISQ-$k$ and AR-ISQ-$k$. We exactly characterize the relevant work of the Sep-ISQ-$k$ system and derive a lower bound on the relevant work of the AR-ISQ-$k$ system.
@@
We characterizing the exact expected relevant work of the Sep-ISQ-$k$ system in Theorems 7.4 and 9.4. As a result, we lower bound the expected relevant work of any $M/G/k$ system under any scheduling policies. In our main results, Theorems 4.1 and 4.2, our Equations 4.1 and 4.3 are based on the Sep-ISQ-$k$ results.
@@
For the AR-ISQ-$k$ system, arbitrary recycling presents multiple difficulties. First, the ISQ-$k$ total work formulas (4.5) and (4.6) cannot be applied directly, so we require specialized test functions, see Section 8.2. Additionally, the changes in the drift caused by the arbitrary recycling events are difficult to characterize exactly. Therefore, we uniformly lower bound the jumps of the test function during these events, allowing us to lower bound the expected relevant work of the AR-ISQ-$k$ system as proved in Theorems 7.6 and 9.6 respectively. We therefore lower bound the expected relevant work of any $M/G/k$ system under arbitrary scheduling policies, resulting in Equations 4.2 and 4.4 in our main results Theorems 4.1 and 4.2. Finally, we can translate lower bounds on the mean relevant work of the $M/G/k$ system into lower bounds on its mean response time under arbitrary scheduling policies.
@@ 5.1. Background on drift method
The drift method [4], also known as the BAR [1] or the rate conservation law [27], states that the average rate of increase and decrease of a random variable must be equal, for any random variable with finite expectation and satisfying certain regularity conditions.
@@
We usually do not directly use the above equation to compute the drift of a test function when we know all the rates at which the system states change. For the ISQ-$k$ system, $w$ increases according to stochastic jumps of size $S$, which arrive according to a Poisson Process with rate $\lambda$. These jumps also cause $i$ to increase by $1/k$ as long as $i<1$. When $i>0$, $w$ decreases at rate $i$ due to work completion. The next lemma is a special case of Section 6.1 in [2], which states that for a piecewise deterministic Markov process, the drift can be characterized by jumps caused by a Poisson process.
@@ 5.2. BAR for unbounded test functions of continuous Markov processes
We say that $g$ belongs to the domain of the generator $G$ of the process $X$ and write $g\in D(G)$ if the above limit exists for all $x\in\mathcal{S}$. The first piece of our novel BAR result is as follows. We prove this result in Appendix A,
@@ 5.3. Applying drift method to ISQ-$k$ for quadratic test functions
If the job size distribution $S$ has a finite third moment, $E[S^{3}]<\infty$, then $\mathbb{E}[W^{2}]<\infty$.
Next, we provide sufficient conditions on the test function $g$ to ensure that $G\circ g$ satisfies Proposition 5.2’s (5.2). This result makes use of the fact that the arrival process is Poisson and we prove in Appendix A.
@@
Suppose $\mathbb{E}[S^{3}]<\infty$ and $g(w,i)=w^{2}+c(w,i)$ is a real-valued function of the ISQ-$k$ system which is twice-differentiable with respect to $w$ for each fixed $i$. Suppose that $|c(w,i)|\leq C_{1}w+C_{2}$ for some constants $C_{1}$ and $C_{2}$ and $\lim_{w\to 0^{+}}g(w,i)=g(0,0)$ for each fixed $i$. Moreover, we assume $|c^{\prime}(w,i)|\leq M_{1}$ and $|c^{\prime\prime}(w,i)|\leq M_{2}$ for some constants $M_{1}$ and $M_{2}$. Then
@@
Now, we are ready to prove our novel BAR result. Combining Lemmas 5.3 and 5.4 we have the following BAR (2.1) result which shows that in the ISQ-$k$ system, for test functions $g$ of the structure described here, the expected value of the $G\circ g$ in steady state is zero. We prove the following lemma in Appendix A.
@@
Suppose $\mathbb{E}[S^{3}]<\infty$ and $g(w,i)=w^{2}+c(w,i)$ is a real-valued function of the ISQ-$k$ system which is twice-differentiable with respect to $w$ for each fixed $i$. Suppose that $|c(w,i)|\leq C_{1}w+C_{2}$ for some constants $C_{1}$ and $C_{2}$ and $\lim_{w\to 0^{+}}g(w,i)=g(0,0)$ for each fixed $i$. Moreover, we assume $|c^{\prime}(w,i)|\leq M_{1}$ and $|c^{\prime\prime}(w,i)|\leq M_{2}$ for some constants $M_{1}$ and $M_{2}$. Then,
@@
We will often apply Lemma 5.5 when the job size distribution follows a truncated distribution such as $S_{x}$ and $S_{\bar{x}}$, for which the assumption that $\mathbb{E}[S^{3}]<\infty$ is automatically satisfied.
It is straightforward to apply Lemma 5.5 to the Sep-ISQ-$k$ system because each subsystem is Markovian and independent. For the AR-ISQ-$k$ system, we only apply Lemma 5.5 to the Poisson arrival stream and deal with the arbitrary recycling stream through a different approach, see [1, 2].
In particular, we specify a version of Lemma 5.1 for the AR-ISQ-$k$ system using the Palm expectation $\mathbb{E}_{r}$ over the moments when jobs recycle, following [1, 2]. Note that the state of the AR-ISQ-$k$ system is $(w,i,a)$, where $w$ is the work, $i$ is the speed, and $a$ is the hidden arrival state. In this paper, we only consider test functions which do not depend on $a$. When it is clear, we write these test functions as $g(w,i)$. For clarity, we write test functions with three inputs $(w,i,a)$ in the following lemma:
@@
For any real-valued differentiable function $g$ of the state of the arbitrary recycling ISQ-$k$ system which does not depend on the hidden arrival state $a$,
$G\circ g(w,i,a)=\lambda_{x}\mathbb{E}[g(w+S,\min\{1,i+1/k\},\cdot)-g(w,i,\cdot)]-\frac{d}{dw}g(w,i,\cdot)i+(\lambda-\lambda_{x})\mathbb{E}_{r\mid w,i,a}[J(W,I)],$
@@
where $J(w,i)=g(w+x,\min\{i+1/k,1\},\cdot)-g(w,i,\cdot)$ denotes the increase in the test function due to the arrival of a size-$x$ job, and $E_{r\mid w,i,a}[\cdot]$ denotes the conditional expectation of recycling with respect to the Palm measure in the immediate future of the state $(w,i,a)$ of the AR-ISQ-$k$ system.
6. Increasing Speed Queue
In this section, we prove that the Sep-ISQ-$k$ system and the AR-ISQ-$k$ system lower bound the relevant work of an $M/G/k$ system under an arbitrary scheduling policy.
Proposition 6.1.
Proof.
We will use the index $j\in\mathcal{J}=\{1,2,...\}$ to denote the busy periods of the ISQ-$k$ system. Specifically, a busy period begins when a job arrives to an empty system, and ends when the system is next empty. Let $W_{j}^{ISQ{\text{-}}k}(t)$ denote the total work in the ISQ-$k$ system consisting of jobs which arrived during busy period $j$ of the ISQ-$k$ system. Note that at all points in time, $W_{j}^{ISQ{\text{-}}k}(t)$ is positive for at most one value of $j$, namely the current busy period. Similarly, let $W_{j}^{k}(t)$ denote the total work remaining in the $k$-server system consisting of jobs which arrived during busy period $i$ of the ISQ-$k$ system.
Thus, we want to show that $B_{j}^{ISQ{\text{-}}k}(t)$ is higher than $B_{j}^{k}(t)$ whenever $W_{j}^{ISQ{\text{-}}k}>0$ , $j$ must be the busy period currently active. Note that $W_{j}^{ISQ{\text{-}}k}(t)=W^{ISQ{\text{-}}k}(t)$, as all of the work in the ISQ-$k$ must have arrived during its current busy period, as the ISQ-$k$ system ends each busy period by completing all work. Thus, $B_{j}^{ISQ{\text{-}}k}(t)$ is $\min\{A_{j}(t)/k,1\}$.
See 4.3
Theorem 6.2.
For an arbitrary job size distribution $S$ and arrival rate $\lambda$, the expected relevant work in the $M/G/k$ system under an arbitrary scheduling policy is lower bounded by the expected relevant work in the separate-recycling-ISQ-$k$ system.
Proof.
We want to lower bound the relevant work in the $M/G/k$. We will divide that work into two categories: Relevant work from jobs with original size $\leq x$, and relevant work from jobs with original size $>x$. We will show that the truncated-ISQ-$k$ system lower bounds the former category, and the M/G/$\infty$ with server speed $1/k$ lower bounds the latter category.
Consider a coupled pair of systems: A truncated $M/G/k$ and the full $M/G/k$. Whenever a job arrives to the full $M/G/k$, if that job has size $\leq x$, a job with the same size arrives to the truncated $M/G/k$. By Poisson splitting, the arrival process to the truncated $M/G/k$ is the desired process.
For jobs with initial size $>x$, the expected relevant work in the separate $M/G/\infty$ system lower bounds the expected relevant work of jobs with initial size $>x$ in the full-$M/G/k$ system because the rate of completion in the $M/G/\infty$ system is always at $1/k$ for each jobs whereas the rate of completion of each job in the full $M/G/k$ is never more than $1/k$.
Thus we see that the expected relevant work in the full-$M/G/k$ system is lower bounded by the expected relevant work in the separate-recycling-ISQ-$k$ system.
We now switch our focus to the arbitrary-recycling ISQ-$k$ system. Recall that we define a recycling to occur when a job of original size greater than $x$ ages down to remaining size of exactly $x$ in the $M/G/k$. From the perspective of relevant work, a recycling event looks like a remaining-size-$x$ job popping into existence, i.e. becoming relevant for the first time. Recyclings happen with rate $\lambda(1-F_{S}(x))$, because every job with original size $>x$ eventually recycles.
In particular, imagine a job of size $x$ arriving into the AR-ISQ-$k$ system whenever a recycling occurs in the $M/G/k$ system. This arrival sequence has a rate $\lambda(1-F_{S}(x))$ and is Markovian.
Theorem 6.3.
For an arbitrary job size distribution $S$ and arrival rate $\lambda$, and an arbitrary $M/G/k$ scheduling policy, there exists a Markovian recycling stream such that the expected relevant work in the $M/G/k$ under the given scheduling policy is lower bounded by the expected relevant work in an arbitrary-recycling-ISQ-$k$ with the given recycling stream.
Proof.
In other words, the relevant work in the full $M/G/k$ matches the total work in a $k$-server system with two arrival streams: Poisson arrivals for jobs with size $\leq x$, and arbitrary arrivals at rate $\lambda(1-F_{S}(x))$ of size exactly $x$, with equivalent scheduling policies.
For this specific arbitrary arrival stream, by Proposition 6.1, we have that the relevant work in the AR-ISQ-$k$ system with the same recycling stream lower bounds the relevant work in the arbitrary $k$-server system and thus lower bounds the relevant work in the full $M/G/k$ system.
Therefore, by Theorem 6.3, the minimum possible mean relevant work in the $M/G/k$ system under an arbitrary scheduling policy is lower bounded by the minimum possible mean work in the AR-ISQ-$k$ system under an arbitrary recycling stream. We lower bound the mean work in the AR-ISQ-$k$ system in Theorem 7.6.

### v3

@@ 4.1. Proof outline
For any job size $S$ and arrival rate $\lambda$, the expected total work in the $M/G/k$ system under any scheduling policy is lower bounded by the expected total work in the ISQ-$k$ system.
We prove Theorem 4.3 using a sample path coupling argument in Section 6.
@@
Using Theorem 4.3, we prove in Theorem 6.2 that the expected relevant work of the Sep-ISQ-$k$ system lower bounds the $M/G/k$ expected relevant work under any scheduling policy for all arrival rates $\lambda$. We also prove in Theorem 6.3 that for any $M/G/k$ scheduling policy there always exists a recycling stream $\{R_{t}\}$ such that expected relevant work of the Rec-ISQ-$k$ system lower bounds that policy’s $M/G/k$ expected relevant work for all arrival rates $\lambda$.
@@
Now that we’ve proven that we can use Sep-ISQ-$k$ and Rec-ISQ-$k$ to lower bound mean relevant work in the M/G/$k$, our remaining goal is to characterize the expected relevant work of the Sep-ISQ-$k$ and Rec-ISQ-$k$ systems. As an intermediate result, a key step is to compute the expected total work of the ISQ-$k$ system. We analyze the expected total work of the ISQ-$k$ system using the DiffeDrift method, see Section 9.1.
@@
Here $v_{1}$ and $u_{1}$ are functions defined as the solutions of differential equations in Definitions 10.1 and 10.2. Equations 4.11 and 4.12 correspond to Propositions 8.3 and 10.3. These results are proved in Sections 8 and 10.
Sep-ISQ-$k$ and Rec-ISQ-$k$ relevant work
@@
Now, it remains to analyze the mean relevant work in Sep-ISQ-$k$ and Rec-ISQ-$k$. We exactly characterize the relevant work of the Sep-ISQ-$k$ system and derive a lower bound on the relevant work of the Rec-ISQ-$k$ system.
@@
We characterize the exact expected relevant work of the Sep-ISQ-$k$ system in Theorems 8.4 and 10.4. As a result, we lower bound the expected relevant work of any $M/G/k$ system under any scheduling policies. In our main results, Theorems 4.1 and 4.2, our Equations 4.4 and 4.9 are based on the Sep-ISQ-$k$ results.
@@
For the Rec-ISQ-$k$ system, recycling presents multiple difficulties. First, the ISQ-$k$ total work formulas (4.11) and (4.12) cannot be applied directly, so we require specialized test functions, see Section 9.2. Additionally, the changes in the drift caused by the recycling events are difficult to characterize exactly. Therefore, we uniformly lower bound the jumps of the test function during these events, allowing us to lower bound the expected relevant work of any Rec-ISQ-$k$ system with any arbitrary recycling stream $\{R_{t}\}$. These results are proved in Theorems 8.6 and 10.6. By Theorem 6.3, we therefore lower bound the expected relevant work of any $M/G/k$ system under arbitrary scheduling policies, resulting in Equations 4.5 and 4.10 in our main results Theorems 4.1 and 4.2.
5. WINE and Existing Lower Bounds
To translate lower bounds on the mean relevant work of the $M/G/k$ system into lower bounds on its mean response time under arbitrary scheduling policies, we use the Work Integral Number Equality (WINE) introduced in Theorem 6.3 of Scully2020. WINE allows us to characterize the mean response time of a generic queueing system under a generic scheduling policy in terms of the mean relevant work for each relevancy-cutoff level $x$.
Proposition 5.1 (WINE Identity).
For an arbitrary scheduling policy $\pi$, in an arbitrary system,
$\mathbb{E}[T^{\pi}]=\dfrac{1}{\lambda}\mathbb{E}[N^{\pi}]=\dfrac{1}{\lambda}\int_{0}^{\infty}\dfrac{\mathbb{E}[W_{x}^{\pi}]}{x^{2}}\,dx.$ (5.1)
WINE has been used to upper bound mean response time under complex scheduling policies in the $M/G/1$ (Scully2022Uniform) as well as in multiserver systems (Grosof2023Optimal and Scully2020).
The first lower bound is given by the resource-pooled $M/G/1$ queue under the single-server optimal SRPT policy. The optimal policy in the resource-pooled $M/G/1$ must lower bound the optimal policy in the $M/G/k$ system, as the resource-pooled model can simulate any $M/G/k$ policy. Since we are using the SRPT relevant work, these are precisely the jobs that SRPT prioritizes. Therefore, the resource-pooled single server SRPT provides a lower bound on the relevant work of any $M/G/k$ system under arbitrary scheduling policy.
The mean relevant work in the SRPT-1 system was characterized by Schrage1966 and is given by:
The second bound is the mean service time bound. To compute the mean relevant work, we imagine jobs arriving into an $M/G/\infty$ queue, whose servers run at the same speed as the $M/G/k$ servers. Each job of initial size $s$ will spend $k\min\{s,x\}$ amount of time as a relevant job, as its size shrinks linearly down to 0. The job thus contribute $\frac{k}{2}\min\{s,x\}^{2}$ amount of relevant work. Therefore,
Taking the maximum of these two bounds for each $x$ and integrating using WINE (Proposition 5.1) gives us our first improved lower bound on the mean response time of the $M/G/k$ under an arbitrary scheduling policy, which we call the MixEx bound, i.e., Equation 4.1 and Equation 4.6.
The primary goal of this paper is to further improve upon MixEx by incorporating the exact mean relevant work of the Sep-ISQ-$k$ system, giving an improved lower bound on mean relevant work in the $M/G/k$, namely Equation 4.9. Combining these bounds with WINE results in a new lower bound which we call the ISQ bound. Finally, the ISQ bound is augmented with a lower bound on the mean relevant work of the Rec-ISQ-$k$ system, giving a final lower bound on the $M/G/k$, namely Equation 4.10, resulting in the final ISQ-Recycling bound.
6. Increasing Speed Queue
In this section, we prove that the Sep-ISQ-$k$ system and the Rec-ISQ-$k$ system lower bound the relevant work of an $M/G/k$ system under an arbitrary scheduling policy.
Proposition 6.1.
Proof.
We will use the index $j\in\mathcal{J}=\{1,2,$...\} to denote the busy periods of the ISQ-$k$ system. Specifically, a busy period begins when a job arrives to an empty system, and ends when the system is next empty. Let $W_{j}^{ISQ{\text{-}}k}(t)$ denote the total work in the ISQ-$k$ system consisting of jobs which arrived during busy period $j$ of the ISQ-$k$ system. Note that at all points in time, $W_{j}^{ISQ{\text{-}}k}(t)$ is positive for at most one value of $j$, namely the current busy period. Similarly, let $W_{j}^{k}(t)$ denote the total work remaining in the $k$-server system consisting of jobs which arrived during busy period $j$ of the ISQ-$k$ system.
Thus, we want to show that $B_{j}^{ISQ{\text{-}}k}(t)$ is higher than $B_{j}^{k}(t)$ whenever $W_{j}^{ISQ{\text{-}}k}>0$. Note that $W_{j}^{ISQ\text{-}k}>0$ only occurs when $j$ is the busy period currently active. Note that $W_{j}^{ISQ{\text{-}}k}(t)=W^{ISQ{\text{-}}k}(t)$ where $j$ is the current busy period at time $t$, as all of the work in the ISQ-$k$ must have arrived during its current busy period, and the ISQ-$k$ system ends each busy period by completing all work. Thus, $B_{j}^{ISQ{\text{-}}k}(t)$ is $\min\{A_{j}(t)/k,1\}$.
See 4.3
Theorem 6.2.
For an arbitrary job size $S$ and arrival rate $\lambda$, the expected relevant work in the $M/G/k$ system under an arbitrary scheduling policy is lower bounded by the expected relevant work in the separate-ISQ-$k$ system.
Proof.
[…]
We prove this theorem using a coupling argument. By coupling, we mean that we construct two different queueing systems on the same probability space and let them evolve in parallel under the exact same arrival process. In particular, every job arrival occurs at the same time in both systems and has the same job size in each system. The only difference between the systems lies in how they process these arrivals.
Consider a coupled pair of systems: a truncated $M/G/k$ system and the full $M/G/k$ system. Whenever a job arrives to the full $M/G/k$, if that job has size $\leq x$, a job with the same size arrives to the truncated $M/G/k$. By Poisson splitting, the arrival process to the truncated $M/G/k$ is the desired process.
For jobs with initial size >x>x, the expected relevant work in the separate $M/G/\infty$ system lower bounds the expected relevant work of jobs with initial size >x>x in the full-$M/G/k$ system because the rate of completion in the $M/G/\infty$ system is always $1/k$ for each job whereas the rate of completion of each job in the full $M/G/k$ is never more than $1/k$.
Thus we see that the expected relevant work in the full-$M/G/k$ system is lower bounded by the expected relevant work in the separate-ISQ-$k$ system.
We now switch our focus to the recycling ISQ-$k$ system. Recall that we define a recycling to occur when a job of original size greater than $x$ ages down to remaining size of exactly $x$ in the $M/G/k$. From the perspective of relevant work, a recycling event looks like a remaining-size-$x$ job popping into existence, i.e. becoming relevant for the first time. Recyclings happen with rate $\lambda(1-F_{S}(x))$, because every job with original size >x>x eventually recycles.
In particular, imagine a job of size $x$ arriving into the Rec-ISQ-$k$ system whenever a recycling occurs in the $M/G/k$ system. This arrival sequence has a rate $\lambda(1-F_{S}(x))$ and is Markovian.
Theorem 6.3.
For an arbitrary job size $S$ and arrival rate $\lambda$, and an arbitrary $M/G/k$ scheduling policy, there exists a Markovian recycling stream $\{R_{t}\}$ such that the expected relevant work in the $M/G/k$ under the given scheduling policy is lower bounded by the expected relevant work in an Rec-ISQ-$k$ with the given recycling stream.
Proof.
In other words, the relevant work in the full $M/G/k$ matches the total work in a $k$-server system with two arrival streams: Poisson arrivals for jobs with size $\leq x$, and recycling arrivals at rate $\lambda(1-F_{S}(x))$ of size exactly $x$, with equivalent scheduling policies.
For this specific arbitrary arrival stream, by Proposition 6.1, we have that the relevant work in the Rec-ISQ-$k$ system with the same recycling stream lower bounds the relevant work in the arbitrary $k$-server system and thus lower bounds the relevant work in the full $M/G/k$ system.
Therefore, by Theorem 6.3, the minimum possible mean relevant work in the $M/G/k$ system under an arbitrary scheduling policy is lower bounded by the minimum possible mean work in the Rec-ISQ-$k$ system under an arbitrary recycling stream. We lower bound the mean work in the Rec-ISQ-$k$ system in Theorem 8.6 for the 2 server case and Theorem 10.6 for the general case.
@@ 7.1. Background on drift method
The drift method (Eryilmaz2012), also known as the BAR (Braverman2017 and Harrison1987) or the rate conservation law (Miyazawa1994), states that the average rate of increase and decrease of a random variable must be equal, for any random variable with finite expectation and satisfying certain regularity conditions.
@@
We usually do not directly use the above equation to compute the drift of a test function when we know all the rates at which the system states change. For the ISQ-$k$ system, $w$ increases according to stochastic jumps of size $S$, which arrive according to a Poisson Process with rate $\lambda$. These jumps also cause $i$ to increase by $1/k$ as long as $i<1$. When $i>0$, $w$ decreases at rate $i$ due to work completion. The next lemma is a special case of Section 6.1 in Braverman2024, which states that for a piecewise deterministic Markov process, the drift can be characterized by jumps caused by a Poisson process.
@@ 7.2. BAR for unbounded test functions of continuous Markov processes
We say that $g$ belongs to the domain of the generator $G$ of the process $X$ and write $g\in D(G)$ if the above limit exists for all $x\in\mathcal{S}$. The first piece of our novel BAR result is as follows. We prove this result in Appendix A.
@@ 7.3. Applying drift method to ISQ-$k$ for quadratic test functions
In an ISQ-$k$ system, if the job size $S$ has a finite third moment, $E[S^{3}]<\infty$, then $\mathbb{E}[W^{2}]<\infty$.
Next, we provide sufficient conditions on the test function $g$ to ensure that $G\circ g$ satisfies Proposition 7.2’s condition. This result makes use of the fact that the arrival process is Poisson and we prove it in Appendix A.
@@
In an ISQ-$k$ system, suppose $\mathbb{E}[S^{3}]<\infty$ and $g(w,i)=w^{2}+c(w,i)$ is a real-valued function of the ISQ-$k$ system which is twice-differentiable with respect to $w$ for each fixed $i$. Suppose that $|c(w,i)|\leq C_{1}w+C_{2}$ for some constants $C_{1}$ and $C_{2}$ and $\lim_{w\to 0^{+}}g(w,i)=g(0,0)$ for each fixed $i$. Moreover, we assume $|c^{\prime}(w,i)|\leq M_{1}$ and $|c^{\prime\prime}(w,i)|\leq M_{2}$ for some constants $M_{1}$ and $M_{2}$. Then
@@
Now, we are ready to prove our novel BAR result. Combining Lemmas 7.3 and 7.4 we have the following BAR (2.1) result which shows that in the ISQ-$k$ system, for test functions $g$ of the structure described here, the expected value of the $G\circ g$ in steady state is zero. We prove the following lemma in Appendix A.
@@
In an ISQ-$k$ system, suppose $\mathbb{E}[S^{3}]<\infty$ and $g(w,i)=w^{2}+c(w,i)$ is a real-valued function of the ISQ-$k$ system which is twice-differentiable with respect to $w$ for each fixed $i$. Suppose that $|c(w,i)|\leq C_{1}w+C_{2}$ for some constants $C_{1}$ and $C_{2}$ and $\lim_{w\to 0^{+}}g(w,i)=g(0,0)$ for each fixed $i$. Moreover, we assume $|c^{\prime}(w,i)|\leq M_{1}$ and $|c^{\prime\prime}(w,i)|\leq M_{2}$ for some constants $M_{1}$ and $M_{2}$. Then,
@@
We will often apply Lemma 7.5 when the job size distribution follows a truncated distribution such as $S_{x}$ and $S_{\bar{x}}$, for which the assumption that $\mathbb{E}[S^{3}]<\infty$ is automatically satisfied. It is straightforward to apply Lemma 7.5 to the Sep-ISQ-$k$ system because each subsystem is Markovian and independent. For the Rec-ISQ-$k$ system, we only apply Lemma 7.5 to the Poisson arrival stream and deal with the recycling stream through a different approach (Braverman2017 and Braverman2024).
In particular, we specify a version of Lemma 7.1 for the Rec-ISQ-$k$ system using the Palm expectation $\mathbb{E}_{\mathfrak{r}}$ over the moments when jobs recycle, following Braverman2017 and Braverman2024. Note that the state of the Rec-ISQ-$k$ system is $(w,i,r)$, where $w$ is the work, $i$ is the speed, and $r$ is the recycling arrival state. In this paper, we only consider test functions which do not depend on $r$. When it is clear, we write these test functions as $g(w,i)$. For clarity, we write test functions with three inputs $(w,i,r)$ in the following lemma:
@@
For any real-valued differentiable function $g$ of the state of the Rec-ISQ-$k$ system which does not depend on the recycling arrival state $r$,
$G\circ g(w,i,r)=\lambda_{x}\mathbb{E}[g(w+S,\min\{1,i+1/k\},\cdot)-g(w,i,\cdot)]-\frac{d}{dw}g(w,i,\cdot)i+(\lambda-\lambda_{x})\mathbb{E}_{\mathfrak{r}\mid w,i,r}[J(W,I)],$
@@
where $J(w,i)=g(w+x,\min\{i+1/k,1\},\cdot)-g(w,i,\cdot)$ denotes the increase in the test function due to the arrival of a size-$x$ job, and $E_{\mathfrak{r}\mid w,i,r}[\cdot]$ denotes the conditional expectation of recycling with respect to the Palm measure in the immediate future of the state $(w,i,r)$ of the Rec-ISQ-$k$ system.